OpenStudy (anonymous):

he molecular formula of table sugar or table sugar (sucrose) is C12H22O11. Reasonably Indicate if 1 mol of sucrose contains: a) 144 g of carbon b) 18 mol carbon atoms c) 5.10 carbon atoms

OpenStudy (frostbite):

I can give you a hint and say the information is in the formula: \[\large n=\frac{ m }{ N }\] Rewrite it: \[\large m=n \times M\] if we got 1 mol then \(n=1 ~ mol\): \[\large m=1 ~ mol \times \to \frac{ m }{ mol }=M\] Based on that can you give me the answer? :)

OpenStudy (frostbite):

I give a little more help: \[\large m(C)=M(C)\] numerical if we forget about units. Say we like \(m(C_{12})\) then we just do this: \[\large m(C_{12})=12 \times M(C)\] \[\large M(C)=12 ~ \frac{ g }{ mol }\]

OpenStudy (frostbite):

Found the answer? :)

OpenStudy (chmvijay):

@Muskan did u get it or noo